# Transfer Learning on Edge Connecting Probability Estimation under Graphon Model

### v1

@@ 2 Preliminary
In this paper, we let $\mathbf{A}\in\{0,1\}^{n\times n}$ denote an adjacency matrix, and $\mathbf{P}$ denote the corresponding probability matrix that generates $\mathbf{A}$, i.e., $\mathbf{A}_{ij}\sim{\sf Ber}(\mathbf{P}_{ij})$. A graphon model typically assumes that each node $i$ has a latent position $u_{i}\in[0,1]$ and $\mathbf{P}_{ij}=f(u_{i},u_{j})$, where $f$ is a symmetric, measurable function on $0,1]^{2}$. Graphon estimation refers to the estimation of the probability matrix $\mathbf{P}$ under the graphon model. The accuracy of the graphon estimation is typically measured using the mean squared error: $\text{MSE}(\hat{\mathbf{P}},\mathbf{P})=\|\mathbf{P}-\hat{\mathbf{P}}\|_{F}^{2}/n^{2}$.
@@ 2.1 Graphon Estimation
Broadly speaking, graphon estimation methods fall into three categories: (1) Global low-rank techniques, including Universal Singular Value Thresholding (chatterjee2015matrix, ; xu2018rates, ); (2) Combinatorial optimization, which directly searches for node assignments (gao2015rate, ; klopp2017oracle, ) (3) Smoothing-based methods, which pool nodes with similar degree (pmlr-v32-chan14, ) or similar latent neighborhoods (zhang2017estimating, ; mukherjee2019graphon, ). Among these methods, neighborhood smoothing (NS) (zhang2017estimating, ) stands out because it can achieve nearly optimal MSE among all estimators computable in polynomial time. The key idea of NS method is to estimate pairwise connection probabilities by measuring the proportion of edges between the respective neighborhoods of node pairs, effectively leveraging local structure to infer global patterns. Further algorithmic details of NS are provided in Appendix E.2.
@@ 2.2 Gromov-Wasserstein Distance
The Gromov-Wasserstein (GW) distance memoli2011gromov ; memoli2014gromov provides a principled framework for comparing metric-measure spaces, making it particularly suitable for comparing graphs of different sizes without requiring node correspondences. GW distance and couplings (and their variations, e.g., fused GW distance vayer2020fused ; nguyen2023linear , sliced GW distance vayer2019sliced ) have proven to be very useful for comparing/aligning graphs, shapes, or distributions supported on different domains chowdhury2019gromov ; arya2024gromov ; xu2019gromov ; gong2022gromov ; maretic2022wasserstein ; xu2019scalable . In its most general formulation, given two cost/distance matrices $C\in{\mathbb{R}}^{m\times m}$ and $D\in{\mathbb{R}}^{n\times n}$, the GW distance between these two distance matrices is defined as: $\min_{\pi\in\Pi(\mu,\nu)}\sum_{i,j,k,\ell}L\big(C(i,k),D(j,l)\big)\,\pi_{ij}\pi_{kl}$ where $\Pi(\mu,\nu)$ is the set of all couplings $\pi\in{\mathbb{R}}^{m\times n}$ such that $\pi_{ij}\geq 0$, $\sum_{ij}\pi_{ij}=1$ and $\sum_{j}\pi_{ij}=\mu_{i}$ and $\sum_{i}\pi_{ij}=\nu_{j}$, and $L$ is the loss function, for example $L(x,y)=(x-y)^{2}$. The minimizer $\pi^{*}$ of the above equation is called an optimal GW coupling. While the GW distance offers a powerful framework, the exact computation of the GW distance is a quadratic assignment problem, which is known to be computationally intensive (memoli2011gromov, ; vogelstein2015fast, ). The GW distance has been widely used for graph alignment and transfer tasks, demonstrating its effectiveness in applications such as node embedding, cross-domain alignment, and subgraph matching (xu2019gromov, ; xu2019scalable, ; chen2020graph, ; kong2023outlier, ; xu2020gromov, ). In our implementation, we compute the GW distance and coupling using the ot.gromov_wasserstein function from the POT library flamary2021pot , which provides a deterministic stationary solution for given inputs and initialization.
Entropic gromov-wasserstein distance for scaling up to large datasets. To alleviate the computational difficulty, solomon2016entropic proposed adding an entropic regularization to the GW objective function as follows: $\min_{\pi\in\Pi(\mu,\nu)}\sum_{i,j,k,\ell}L\big(C(i,k),D(j,l)\big)\,\pi_{ij}\pi_{kl}+{\epsilon}{\sf KL}\left(\pi\mid\mu\otimes\nu\right)\,.$ Although the addition of an entropic regularizer does not make the optimization problem convex (unless $\epsilon$} exceeds a certain threshold), it significantly enhances computational efficiency. As elaborated in solomon2016entropic , this regularization enables a Sinkhorn-like algorithm, adapted from cuturi2013sinkhorn for the standard optimal transport problem, which is simple to implement and typically exhibits fast convergence rioux2024entropic .
@@ 3.2 Proposed Method
Overview. Figure 1 shows the workflow of our method GTrans, which consists of three main steps: (1) Initial graphon estimation: GTrans begins with observed source and target networks and computes their respective initial estimates $\hat{\mathbf{P}}^{ini}_{s}\in[0,1]^{n_{s}\times n_{s}}$, $\hat{\mathbf{P}}^{ini}_{t}\in[0,1]^{n_{t}\times n_{t}}$. We obtain these initial edge probability matrices using the neighborhood smoothing (NS) method zhang2017estimating , which provides consistent and robust estimates that serve as the input for the subsequent GW alignment and debiasing steps. These initial estimators capture the basic structure of the respective graphs but may have limited accuracy for the target graph due to its smaller size. (2) Transferring step: We employ optimal transport to calculate the alignment matrix between $\pi$ between the source and target domains. Using this alignment, we transfer the source estimate to the target domain to obtain $\hat{\mathbf{P}}_{t}^{\text{trans}}$, which is further smoothed to obtain $\hat{\mathbf{P}}_{t}^{\text{trans2}}$. If the source-target domain shift (measured by the transport distance $d$) is below a threshold $\delta$, our final estimate is $\hat{\mathbf{P}}^{t}=\hat{\mathbf{P}}_{t}^{\text{trans2}}$. (3) Debiasing step: When $d>\delta$, indicating a significant domain shift, direct transfer may miss target-specific structures. To correct this, we perform a debiasing step. We compute the residual matrix $\mathbf{R}_{t}=\hat{\mathbf{P}}^{ini}_{t}-\hat{\mathbf{P}}_{t}^{\text{trans2}}$, which captures structural patterns present in the target graph but not explained by transfer from the source. However, $\mathbf{R}_{t}$ is noisy due to the small target sample. To extract a meaningful signal, we apply neighborhood smoothing on $\mathbf{R}_{t}$ to obtain a denoised residual estimate $\hat{\mathbf{P}}^{\text{res}}$. This smoothed correction is then added back to the transferred estimate, yielding the final output $\hat{\mathbf{P}}^{t}=\hat{\mathbf{P}}_{t}^{\text{trans2}}+\hat{\mathbf{P}}^{\text{res}}$.
@@ 3.2.2 Debiasing Step
When source and target graphons differ significantly, negative transfer can occur. We quantify this domain difference using the GW between $\hat{\mathbf{P}}_{t}^{ini}$ and $\hat{\mathbf{P}}_{s}^{ini}$. When this distance exceeds a predetermined threshold, we implement our debiasing procedure to mitigate potential negative transfer effects. Specifically, we compute a residual matrix: $\mathbf{R}_{t}=\hat{\mathbf{P}}_{t}^{ini}-\hat{\mathbf{P}}_{t}^{trans2}$. This residual represents the unexplained structural information in the target graph, i.e., the patterns that weren’t captured by the transferred source graphon estimator. We apply neighborhood smoothing to this residual to obtain a smoothed residual estimator: $\hat{\mathbf{P}}_{t}^{res}$. This step retains meaningful target-specific structures while reducing noise, adaptively correcting the transferred estimator and avoiding degeneration to the target-only NS baseline. Finally, we combine the transferred estimator with the smoothed residual to get our final graphon estimator for the target graph: $\hat{\mathbf{P}}_{t}=\hat{\mathbf{P}}_{t}^{trans2}+\hat{\mathbf{P}}_{t}^{res}$. However, if the distance is smaller than the threshold, we simply use the smoothed transferred estimator $\hat{\mathbf{P}}_{t}^{trans2}$ as our final estimator. Algorithm 1 shows our complete procedure, where the optimal transport can be calculated using either GW or its entropic variant EGW, and the choice depends primarily on the computational considerations. Although transfer works best when the source and target graphs are structurally similar, GTrans remains stable even when they differ. In practice, we measure similarity using the GW distance between the initial estimators $\hat{\mathbf{P}}_{s}^{\text{ini}}$ and $\hat{\mathbf{P}}_{t}^{\text{ini}}$. If this distance is large, the debiasing step allows GTrans to perform on par with the target-only NS method, ensuring reliability even without strong alignment.
@@ 4 Theoretical Properties
In this section, we present a stability analysis of the estimated alignment matrix obtained via minimizing the entropic Gromov-Wasserstein distance.
Although EGW has been successfully applied in a variety of domains wang2023neural ; klein2024genot , the question of the stability of the optimizer $\pi^{*}$ with respect to perturbations in the cost or distance matrices $(C,D)$ has received limited attention. Very recently, zhang2024gromov first established duality theory for EGW, deriving optimal $n^{-1/2}$ empirical convergence rates and proving stability with respect to the regularization parameter and rioux2024entropic analyzed the a certain aspect of stability of the estimator, when both $C$ and $D$ are euclidean distance matrices based on $m$ and $n$ observations respectively. These analyses do not address the case when (i) $C$ and $D$ are arbitrary cost matrices without the inner-product structure of Euclidean distances, and (ii) the stability of the EGW estimator under perturbations of the cost functions, both of which are particularly relevant to our transfer learning framework. Recall that in the GTrans algorithm, the estimated alignment matrix $\hat{\pi}$ is obtained using EGW with inputs $C=\hat{\mathbf{P}}_{s}^{\rm ini}$ and $D=\hat{\mathbf{P}}_{s}^{\rm ini}$. As a result, these matrices cannot be interpreted as Euclidean distance matrices between point clouds, since they do not arise from pairwise distances in an inner-product space. It is therefore natural to ask how far $\hat{\pi}$ is from the oracle alignment matrix $\pi^{*}$, which minimizes the same objective using the true connectivity matrices $C=\mathbf{P}_{s}$ and $D=\mathbf{P}_{t}$. While the stability of convex optimization problems has been extensively studied (see, e.g., bonnans2013perturbation for an overview of techniques and results), these tools do not directly apply to our setting due to the nonconvex nature of the EGW objective. Our subsequent theorem showcases that the distance between $\hat{\pi}$ and $\pi^{*}$ can be upper bounded by a distance between $\hat{\mathbf{P}}^{\rm ini}_{s}$ and $\mathbf{P}_{s}$, and $\hat{\mathbf{P}}^{\rm ini}_{t}$ and $\mathbf{P}_{t}$:

### v2

@@ 2 Preliminary
In this paper, we let $\mathbf{A}\in\{0,1\}^{n\times n}$ denote an adjacency matrix, and $\mathbf{P}$ denote the corresponding probability matrix that generates $\mathbf{A}$, i.e., $\mathbf{A}_{ij}\sim{\sf Ber}(\mathbf{P}_{ij})$. A graphon model typically assumes that each node $i$ has a latent position $u_{i}\in[0,1]$ and $\mathbf{P}_{ij}=f(u_{i},u_{j})$, where $f$ is a symmetric, measurable function on $0,1]^{2}$. Graphon estimation refers to the estimation of the probability matrix $\mathbf{P}$ under the graphon model. The accuracy of the graphon estimation is typically measured using the mean squared error: $\text{MSE}(\hat{\mathbf{P}},\mathbf{P})=\|\mathbf{P}-\hat{\mathbf{P}}\|_{F}^{2}/n^{2}$. MSE is widely used as a standard evaluation metric in existing literature (zhang2017estimating, ; qin_iterative_2021, ; pmlr-v32-chan14, ).
@@ 2.1 Graphon Estimation
Broadly speaking, graphon estimation methods fall into three categories: (1) Global low-rank techniques, including Universal Singular Value Thresholding (chatterjee2015matrix, ; xu2018rates, ); (2) Combinatorial optimization, which directly searches for node assignments (gao2015rate, ; klopp2017oracle, ) (3) Smoothing-based methods, which pool nodes with similar degree (pmlr-v32-chan14, ) or similar latent neighborhoods (zhang2017estimating, ; mukherjee2019graphon, ). Among these methods, neighborhood smoothing (NS) (zhang2017estimating, ) stands out because it can achieve nearly optimal MSE among all estimators computable in polynomial time. The key idea of NS method is to estimate pairwise connection probabilities by measuring the proportion of edges between the respective neighborhoods of node pairs, effectively leveraging local structure to infer global patterns. Further algorithmic details of NS are provided in Appendix G.2.
@@ 2.2 Gromov-Wasserstein Distance
The Gromov-Wasserstein (GW) distance memoli2011gromov ; memoli2014gromov provides a principled framework for comparing metric-measure spaces, making it particularly suitable for comparing graphs of different sizes without requiring node correspondences. GW distance and couplings (and their variations, e.g., fused GW distance vayer2020fused ; nguyen2023linear , sliced GW distance vayer2019sliced ) have proven to be very useful for comparing/aligning graphs, shapes, or distributions supported on different domains chowdhury2019gromov ; arya2024gromov ; xu2019gromov ; gong2022gromov ; maretic2022wasserstein ; xu2019scalable . In its most general formulation, given two cost/distance matrices $C\in{\mathbb{R}}^{m\times m}$ and $D\in{\mathbb{R}}^{n\times n}$, the GW distance between these two distance matrices is defined as: $\min_{\pi\in\Pi(\mu,\nu)}\sum_{i,j,k,\ell}L\big(C(i,k),D(j,l)\big)\,\pi_{ij}\pi_{kl}$ where $\Pi(\mu,\nu)$ is the set of all couplings $\pi\in{\mathbb{R}}^{m\times n}$ such that $\pi_{ij}\geq 0$, $\sum_{ij}\pi_{ij}=1$ and $\sum_{j}\pi_{ij}=\mu_{i}$ and $\sum_{i}\pi_{ij}=\nu_{j}$, and $L$ is the loss function, for example $L(x,y)=(x-y)^{2}$.Here, $\mu$ and $\nu$ denote the uniform measures on the source and target nodes, respectively; that is, $\mu=(1/n_{S},\ldots,1/n_{S})$ and $\nu=(1/n_{T},\ldots,1/n_{T})$. The minimizer $\pi^{*}$ of the above equation is called an optimal GW coupling. While the GW distance offers a powerful framework, the exact computation of the GW distance is a quadratic assignment problem, which is known to be computationally intensive (memoli2011gromov, ; vogelstein2015fast, ). The GW distance has been widely used for graph alignment and transfer tasks, demonstrating its effectiveness in applications such as node embedding, cross-domain alignment, and subgraph matching (xu2019gromov, ; xu2019scalable, ; chen2020graph, ; kong2023outlier, ; xu2020gromov, ).
Entropic gromov-wasserstein distance for scaling up to large datasets. To alleviate the computational difficulty, solomon2016entropic proposed adding an entropic regularization to the GW objective function as follows: $\min_{\pi\in\Pi(\mu,\nu)}\sum_{i,j,k,\ell}L\big(C(i,k),D(j,l)\big)\,\pi_{ij}\pi_{kl}+{\epsilon}{\sf KL}\left(\pi\mid\mu\otimes\nu\right)\,.$ The product $\mu\otimes\nu$ is the product measure assigning a uniform weight of $(n_{S}n_{T})^{-1}$. Although the addition of an entropic regularizer does not make the optimization problem convex (unless $\epsilon$} exceeds a certain threshold), it significantly enhances computational efficiency. As elaborated in solomon2016entropic , this regularization enables a Sinkhorn-like algorithm, adapted from cuturi2013sinkhorn for the standard optimal transport problem, which is simple to implement and typically exhibits fast convergence rioux2024entropic .
@@ 3.2 Proposed Method
Overview. Figure 1 shows the workflow of our method GTrans, which consists of three main steps: (1) Initial graphon estimation: GTrans begins with observed source and target networks and computes their respective initial estimates $\hat{\mathbf{P}}^{ini}_{s}\in[0,1]^{n_{s}\times n_{s}}$, $\hat{\mathbf{P}}^{ini}_{t}\in[0,1]^{n_{t}\times n_{t}}$. We obtain these initial edge probability matrices using the NS method zhang2017estimating . These initial estimators capture the basic structure of the respective graphs but may have limited accuracy for the target graph due to its smaller size. (2) Transferring step: We employ optimal transport to calculate the alignment matrix between $\pi$ between the source and target domains. Using this alignment, we transfer the source estimate to the target domain to obtain $\hat{\mathbf{P}}_{t}^{\text{trans}}$, which is further smoothed to obtain $\hat{\mathbf{P}}_{t}^{\text{trans2}}$. If the source-target domain shift (measured by the transport distance $d$) is below a threshold $\delta$, our final estimate is $\hat{\mathbf{P}}^{t}=\hat{\mathbf{P}}_{t}^{\text{trans2}}$. (3) Debiasing step: When $d>\delta$, indicating a significant domain shift, direct transfer may miss target-specific structures. To correct this, we perform a debiasing step. We compute the residual matrix $\mathbf{R}_{t}=\hat{\mathbf{P}}^{ini}_{t}-\hat{\mathbf{P}}_{t}^{\text{trans2}}$, which captures structural patterns present in the target graph but not explained by transfer from the source. However, $\mathbf{R}_{t}$ is noisy due to the small target sample. To extract a meaningful signal, we apply neighborhood smoothing on $\mathbf{R}_{t}$ to obtain a denoised residual estimate $\hat{\mathbf{P}}^{\text{res}}$. This smoothed correction is then added back to the transferred estimate, yielding the final output $\hat{\mathbf{P}}^{t}=\hat{\mathbf{P}}_{t}^{\text{trans2}}+\hat{\mathbf{P}}^{\text{res}}$.
@@ 3.2.2 Debiasing Step
When source and target graphons differ significantly, negative transfer can occur. We quantify this domain difference using the GW between $\hat{\mathbf{P}}_{t}^{ini}$ and $\hat{\mathbf{P}}_{s}^{ini}$. When the distance is smaller than the threshold ($d<\delta$), we simply use the smoothed transferred estimator $\hat{\mathbf{P}}_{t}^{trans2}$ as our final estimator. When this distance exceeds a predetermined threshold ($d>\delta$), we implement the following debiasing procedure to mitigate potential negative transfer effects.
Specifically, we first compute a residual matrix: $\mathbf{R}_{t}=\hat{\mathbf{P}}_{t}^{ini}-\hat{\mathbf{P}}_{t}^{trans2}$. This residual matrix captures (a) target-specific structural patterns not explained by the transferred estimator and (b) random noise due to small sample size of the target graph. To keep the meaningful structure from part (a) but remove the noise from part (b), we then use the neighborhood smoothing method (zhang2017estimating, ). This method averages information from nearby nodes with similar connection patterns, which smooths out random fluctuations and keeps the stable, informative patterns. The result is a cleaner, smoothed residual estimator, $\hat{\mathbf{P}}_{t}^{res}$. Finally, we combine the transferred estimator with the smoothed residual to get our final graphon estimator for the target graph: $\hat{\mathbf{P}}_{t}=\hat{\mathbf{P}}_{t}^{trans2}+\hat{\mathbf{P}}_{t}^{res}$. Adding the smoothed residual back allows the model to recover target-specific structural information that was missing from the transferred estimator, without reintroducing the random noise filtered out during smoothing.
Algorithm 1 summarizes our complete procedure, where the optimal transport can be calculated using either GW or its entropic variant EGW, and the choice depends primarily on the computational considerations.
@@ 4 Theoretical Properties
In this section, we present a stability analysis of the estimated alignment matrix obtained via minimizing the entropic Gromov-Wasserstein distance. Although EGW has been successfully applied in a variety of domains wang2023neural ; klein2024genot , the question of the stability of the optimizer $\pi^{*}$ with respect to perturbations in the cost or distance matrices $(C,D)$ has received limited attention. Very recently, zhang2024gromov first established duality theory for EGW, deriving optimal $n^{-1/2}$ empirical convergence rates and proving stability with respect to the regularization parameter and rioux2024entropic analyzed the a certain aspect of stability of the estimator, when both $C$ and $D$ are euclidean distance matrices based on $m$ and $n$ observations respectively. These analyses do not address the case when (i) $C$ and $D$ are arbitrary cost matrices without the inner-product structure of Euclidean distances, and (ii) the stability of the EGW estimator under perturbations of the cost functions, both of which are particularly relevant to our transfer learning framework. Recall that in the GTrans algorithm, the estimated alignment matrix $\hat{\pi}$ is obtained using EGW with inputs $C=\hat{\mathbf{P}}_{s}^{\rm ini}$ and $D=\hat{\mathbf{P}}_{s}^{\rm ini}$. As a result, these matrices cannot be interpreted as Euclidean distance matrices between point clouds, since they do not arise from pairwise distances in an inner-product space. It is therefore natural to ask how far $\hat{\pi}$ is from the oracle alignment matrix $\pi^{*}$, which minimizes the same objective using the true connectivity matrices $C=\mathbf{P}_{s}$ and $D=\mathbf{P}_{t}$. While the stability of convex optimization problems has been extensively studied (see, e.g., bonnans2013perturbation for an overview of techniques and results), these tools do not directly apply to our setting due to the nonconvex nature of the EGW objective. Our subsequent theorem showcases that the distance between $\hat{\pi}$ and $\pi^{*}$ can be upper bounded by a distance between $\hat{\mathbf{P}}^{\rm ini}_{s}$ and $\mathbf{P}_{s}$, and $\hat{\mathbf{P}}^{\rm ini}_{t}$ and $\mathbf{P}_{t}$:
@@
The following result shows that the empirical GW distance between initial estimators $\hat{\mathbf{P}}_{S}^{\mathrm{init}}$ and $\hat{\mathbf{P}}_{T}^{\mathrm{init}}$ closely approximates the population GW distance between $\mathbf{P}_{S}$ and $\mathbf{P}_{T}$, up to small estimation error.
Theorem 4.2.
Let $\delta_{n_{S}}$ and $\delta_{n,T}$ denote the estimation error of $\mathbb{P}_{S}$ and $\mathbb{P}_{T}$ respectively with respect to average squared Frobenious norm, i.e.
$\mathbb{P}\left(\frac{1}{n_{S}^{2}}\|\hat{\mathbf{P}}^{\rm init}_{S}-\mathbf{P}_{S}\|_{F}^{2}\leq\delta_{n_{S}}\right)\geq 1-{\epsilon}_{n},\quad\mathbb{P}\left(\frac{1}{n_{T}^{2}}\|\hat{\mathbf{P}}^{\rm init}_{T}-\mathbf{P}_{T}\|_{F}^{2}\leq\delta_{n_{T}}\right)\geq 1-{\epsilon}_{n}\,.$
Define $\delta_{n}=\delta_{n_{S}}+\delta_{n_{T}}$. Then, on the intersection of the above events, we have:
$\mathrm{GW}^{2}_{2}(\mathbf{P}_{s},\mathbf{P}_{t})$ $\leq 4\left(\mathrm{GW}^{2}_{2}(\hat{\mathbf{P}}^{\rm init}_{S},\hat{\mathbf{P}}^{\rm init}_{T})+\delta_{n}\right)$
$\mathrm{GW}^{2}_{2}(\hat{\mathbf{P}}^{\rm init}_{S},\hat{\mathbf{P}}^{\rm init}_{T})$ $\leq 4\left(\mathrm{GW}^{2}_{2}(\mathbf{P}_{s},\mathbf{P}_{t})+\delta_{n}\right)\,.$
We next justify the theoretical validity of the alignment step. The following result shows that applying the optimal GW coupling preserves $L_{2}$ proximity between the source and target graphons.
Theorem 4.3.
Let $\pi^{*}$ be the optimal coupling of $(f,g)$ for GW-distance, i.e.
$\pi^{*}={\arg\min}_{\pi\in\Gamma}\int\int(f(x,x^{\prime})-g(y,y^{\prime}))^{2}\ d\pi(x,y)\ d\pi(x^{\prime},y^{\prime})\,.$
Define $\tilde{g}$ (population analogue of $\hat{\pi}^{\top}\hat{P}_{s}\hat{\pi}$) as:
$\tilde{g}(y,y^{\prime})=\int_{[0,1]^{2}}\pi^{*}(y,x)f(x,x^{\prime})\pi^{*}(x^{\prime},y)\ dx\ dx^{\prime}\,.$
Then we have,
$\|\tilde{g}-g\|_{2}^{2}\leq GW_{2}^{2}(f,g)\,.$
As a consequence, when $\mathrm{GW}_{2}(f,g)$ is small, then $\tilde{g}$ is close to $g$, which validates the alignment strategy.
Remark 4.4.
The estimation error $\delta_{n_{S}},\delta_{n_{T}}$ depends on the estimation strategy and the underlying data-generating process. For example, of $\mathbf{P}_{s}$ and $\mathbf{P}_{T}$ are generated from a $\alpha$-Hölder graphon, then it is possible to construct $\hat{\mathbf{P}}_{s}$ and $\hat{\mathbf{P}}_{T}$ such that $\delta_{n}\asymp n^{-2\alpha/(\alpha+1)}$ for $0<\alpha<1$ and $\delta_{n}\asymp(\log n)/n$ for $\alpha\geq 1$ (e.g., see gao2015rate ) for $n\in\{n_{S},n_{T}\}$, albeit they are NP-hard to compute. Later, zhang2017estimating proposed a neighborhood smoothing based approach to obtain $(\hat{\mathbf{P}}_{S},\hat{\mathbf{P}}_{t})$ with $\delta_{n}\asymp\sqrt{\log{n}/n}$ for $n\in\{n_{S},n_{T}\}$.